Mrs. Smith needed 2 3/4 cups of sugar to make fudge and 2 3/8 cups to make ice cream. How much did she need all together?
Answer by stanbon(75887)

Mrs. Smith needed 2 3/4 cups of sugar to make fudge and 2 3/8 cups to make ice cream. How much did she need all together?
-----------------------------------------
2 3/4 + 2 3/8
= 11/4 + 19/8
= 22/8 + 19/8
= 41/8
= 5 1/8 cups
===================
Cheers,
Stan H.